# Hierarchical Refinement of Universal Multimodal Attacks on Vision-Language Models

### v1

@@ Abstract
To overcome this limitation, we propose Hierarchical Refinement Attack (HRA), a multimodal universal attack framework for VLP models. HRA refines universal adversarial perturbations (UAPs) at both the sample level and the optimization level, improving transferability to effectively break cross-modal alignment across models, datasets and tasks. For the image modality, we disentangle adversarial examples into clean images and perturbations, allowing each component to be handled independently for more effective disruption of cross-modal alignment. We further introduce a ScMix augmentation strategy that diversifies visual contexts and strengthens both global and local utility of UAPs, thereby reducing reliance on spurious features. In addition, we refine the optimization path by leveraging a temporal hierarchy of historical and estimated future gradients to avoid local minima and stabilize universal perturbation learning. For the text modality, HRA identifies globally influential words by combining intra-sentence and inter-sentence importance measures, and subsequently utilizes these words as universal text perturbations. Extensive experiments across various downstream tasks, VLP models, and datasets, including cross-modal retrieval scenarios, visual grounding and other downstream tasks, demonstrate the superiority of the proposed universal multimodal attacks.
@@
Vision-Language Pre-trained (VLP) models play a crucial role in bridging the gap between images and texts, enabling multimodal understanding, generation, and response. These models take advantage of abundant information from large-scale unlabelled datasets via self-supervised pre-training, e.g., CLIP [29], ALBEF [16], and TCL [39]. The knowledge acquired during the pre-training phase can be effectively transferred to various downstream tasks with minimal additional training, e.g., fine-tuning. This transferability makes VLP models highly versatile and efficient, allowing them to adapt to new tasks and datasets with relatively low computational costs. Due to their promising performance, VLP models have been widely applied to various vision-language tasks, ranging from multimodal retrieval [22] to image captioning [54].
@@
As the deployment of VLP models continues to grow, evaluating their robustness has become essential. This is vitally important for real-world applications, especially in sensitive and high-stakes applications, where potential threats and uncertainties demand models that are both robust and reliable. Furthermore, uncovering model vulnerabilities would drive advancements in the development of more resilient and reliable neural networks. Adversarial attacks serve as one of the most critical strategies for assessing model vulnerabilities and have been widely studied and applied [31, 50, 25, 3, 23, 48, 43]. These attacks primarily involve crafting imperceptible perturbations that, when injected into data, can mislead models into making incorrect predictions.
@@
Adversarial attacks were originally designed for unimodal scenarios, where only a single data modality is considered [31, 49, 25]. Recently, several studies have begun exploring specific multimodal tasks, e.g., image-text retrieval [44, 52], and visual question answering [40]. These methods typically rely on task-specific losses or model outputs and thus are not directly applicable to VLP models. For example, Zhang et al. [44, 52] propose to learn adversarial perturbations by enlarging the similarity between clean and adversarial data guided by a similarity matrix. Yin et al. [40] introduce a masked answer anti-recovery to prevent models from recovering the correct answer tokens for each masked text. Current methods for VLP attacks mainly concentrate on features that are independent of specific tasks and output, i.e., enlarging the feature distance between adversarial data and original data to break the cross-modal alignments [42]. Further, Lu et al. [24, 43] increase data diversity to better exploit cross-modal interactions, thereby promoting the transferability of perturbations. However, these methods customize adversarial perturbations for each sample based on specific characteristics of each sample. Sample-wise approaches have to learn adversarial perturbations for new data from scratch, resulting in significant computational overhead. This would significantly deter their applications from large-scale cases. Zhou et al. [55, 47] make early attempts to learn universal adversarial attacks for the image modality. However, the method in [55] is restricted to specific models, and neither work exploits text-based attacks, which could further enhance adversarial effectiveness. Fang et al. [11] consider both image and text universal attacks, where the text-side attacks are implemented through word substitutions. These substitutions are obtained by first learning adversarial embeddings and then searching the corpus for words that are semantically aligned with these embeddings. However, this method requires access to a predefined word library, and the mismatch between the learned adversarial embeddings and their final token-level realizations can weaken the attack effectiveness. In addition, these methods do not consider the characteristics of different tasks, thus limiting the transferability.
To enable effective and transferable adversarial attacks, we first identify the fundamental obstacles. First, different VLP models are pre-trained with diverse architectures, learning objectives, and datasets, and are further fine-tuned for specific downstream tasks [29, 16, 39]. These variations lead different VLP models to rely on different features or patterns of images to make decisions, making it difficult to craft transferable adversarial perturbations. For instance, image-text retrieval returns relevant data from one modality for a query from the other by ranking the global image-text similarity between them, while image captioning generates descriptive text for an image by recognizing objects, understanding scenes, and capturing relationships. Thus, attacking retrieval models requires disrupting ranking orders based on global similarity, whereas attacking captioning models demands precise semantic manipulation to mislead content generation. Second, the heterogeneity between data adds another layer of complexity. Images and texts inherently differ in format and describe the same underlying content from distinct perspectives, each carrying diverse and complementary information. This causes the complex cross-modal interactions, making UAPs hardly maximally disrupt the learned cross-modal alignment. Finally, the data-independent nature of UAPs further makes it harder to achieve both effectiveness and transferability. Unlike sample-wise perturbations, UAPs are learned on the entire dataset, which cannot make use of specific characteristics and vulnerabilities of each sample. Since each sample follows its own patterns, UAPs have to be strong enough to break diverse patterns for effective attacks. As a result, the learning process often leads to shortcut exploitation, where UAPs overfit to specific model weaknesses, making them effective against source models (the attacking target during training) but ultimately causing overfitting and reducing transferability. In summary, understanding and manipulating complex cross-modal relationships in sample-, modality-, and task-agnostic settings is particularly challenging, hindering the development of effective and transferable attacks.
@@
To overcome these obstacles, we propose Hierarchical Refinement Attack (HRA). HRA enhances the universal adversarial transferability by focusing on sample-level and optimization-level refinements to disrupt cross-modal alignments. First, HRA disentangles adversarial examples, treating original data and UAPs separately to better exploit cross-modal interactions. Specifically, a novel semantic-preserving data augmentation technique called ScMix is designed to increase data diversity and facilitate the exploration of cross-modal interactions. At the same time, we enhance both the global and local utilities of UAPs, which can avoid the generation of UAPs relying on shortcuts to surrogate models. Additionally, we investigate the optimization trajectory, which is closely related to overfitting. Overfitting often occurs when the optimization process of gradient-based methods converges to local optima. To mitigate this problem, we propose future-aware momentum, which incorporates historical gradients and predicted future gradients to regularize the current update and prevent the optimization from local optima. Furthermore, we implement a simple yet effective adversarial text attack that leverages intra- and inter-sentence importance measures to extract influential words as uniform adversarial perturbations. Through extensive experiments, we demonstrate the effectiveness of the proposed method.
The main contributions of this work are summarized as follows:
This paper identifies the key characteristics and challenges of universal attacks on VLP models, including the sample-agnostic, model-agnostic, and task-agnostic nature of UAPs, as well as the insufficient exploration of cross-modal interactions.
A novel HRA method is proposed, which learn multi-modal universal adversarial perturbations. HRA performs hierarchical refinement of adversarial examples by disentangling adversarial data and making use of both past and future gradients to exploit intrinsic cross-modal interactions and regularize the optimization path, alleviating overfitting.
@@
Thorough experiments are conducted, including a wide range of target VLP models, downstream tasks and datasets, including cross-modal retrieval scenarios, visual grounding and other downstream tasks. The results demonstrate that the proposed method effectively improves the transferability of universal multimodal attacks.
Compared to our previous conference work, i.e., ETU [47], our HRA introduces several substantial improvements. First, we conduct a deeper and more comprehensive investigation into the challenges of achieving transferable universal attacks. Second, we design novel strategies to enhance transferability, i.e., future-aware Momentum. We compute both the previous and future gradients, which are closely related to overfitting. The future-aware momentum strategy effectively mitigates overfitting, resulting in more transferable attacks. Third, we expand our evaluation by comparing our method against a broader range of state-of-the-art approaches and testing it across more target and source models. We also provide more in-depth analysis to offer deeper insights into the task and further highlight the superiority of our proposed method.
@@
The rest of the paper is organized as follows: Section II provides a brief review of related works. Section III presents a detailed introduction to our methodology. In Section IV, we conduct comprehensive experiments, including attacks on various VLP models, datasets, and downstream tasks, as well as ablation studies and visualizations, to demonstrate the effectiveness of the proposed method. Finally, Section V concludes the paper and discusses current limitations and potential directions for future research.
@@ II-A Vision-Language Pre-training
Digital and realistic worlds encompass diverse data types, e.g., image, text, audio, video, and so on. To better understand the world, vision-language models (VLMs) are designed to understand visual and language content, capture their relations and generate multimodal content. These models facilitate diverse tasks e.g., image-text retrieval [51, 53, 46], image captioning [6], visual question answering [1], visual entailment [37], and so on [7, 29, 45]. VLMs typically require a large number of high-quality training data to establish effective prediction ability, which is time-consuming and restricted by the availability of large-scale data. To address these obstacles, vision-language pre-training (VLP) has been developed. It harnesses vast amounts of unlabeled multi-modal data, using self-supervised learning techniques to gain representative feature learning ability [15, 29, 16, 39]. VLP models exhibit outstanding generalizability and transferability, allowing adaptation to diverse datasets and downstream tasks through fine-tuning, thereby offering cost-effectiveness and flexibility. Representative VLP methods include CLIP [29], BLIP [15], ALBEF [16], and TCL [39]. CLIP aligns visual-linguistic modalities by projecting images and texts into a unified feature space, using contrastive learning to maximize similarity for matched pairs and increase the distance for unmatched pairs. BLIP [15] enhances dataset quality by synthesizing captions for web images, employing three contrastive learning objectives to jointly pre-train the model, i.e., image-text contrastive learning, image-text matching, and image-conditioned language modelling. ALBEF [16] first aligns unimodal representations of pair images and texts, and then fuses them with cross-modal attention to obtain joint representations. TCL [39] utilizes contrastive learning to perform both inter- and intra-modal alignment while preserving mutual information between global and local representations in each modality.
@@ II-B Adversarial Attack
Real-world applications often involve uncertainty, randomness, and potential threats, making it crucial to ensure that developed models are both robust and secure. Adversarial attacks are one of the most effective methods for evaluating model robustness, which aims to induce incorrect predictions of target models by injecting imperceptible perturbations (i.e., adversarial perturbations) into input data [31, 44, 52, 50]. According to the scope, existing adversarial perturbations can be roughly categorized into sample-wise [25, 3], class-wise [2] and universal adversarial perturbations (UAPs) [26, 17, 20], which are designed for individual samples, specific classes, and all data, respectively. Generally, sample-wise and class-wise perturbations are more effective than universal perturbations as they can exploit vulnerabilities specific to each sample or class. However, they are less transferable to unseen data and must be generated from scratch for new samples or classes, limiting their scalability in large-scale applications. In contrast, universal perturbations are more efficient, as they can be reused across different data once learned. Nevertheless, since UAPs are independent of specific data, they fail to leverage data-specific characteristics and are more prone to overfit to target models during training, limiting their effectiveness and transferability.
@@
Early methods learn adversarial perturbations in white-box settings, assuming access to information about the victim models, tasks, and data [31, 44, 52]. However, in real-world applications, such target information might not be available and it is impractical to generate perturbations for all possible models, tasks, and datasets. Consequently, adversarial perturbations must often be learned in the black-box settings where target information is inaccessible, and the perturbations are required to be transferable to different models, tasks and data. However, due to factors like architectural differences, task variations and data distribution shifts, white-box attack methods would lose efficacy when applied to black-box scenarios. To address this challenge, current solutions primarily fall into two categories: query-based attacks [27, 5, 56] and transfer-based attacks [21, 9, 35]. Query-based attacks learn adversarial perturbations by analyzing models’ feedback on perturbed inputs. As the inner architecture is not accessible, these attacks often estimate gradients using techniques like the finite difference approximation [5]. However, query-based attacks still require the existence of target models, limiting their practicality. Transfer-based methods overcome this limitation by leveraging substitute models, assuming that adversarial perturbations generated for one model may transfer to others. However, adversarial transferability is not always reliable due to differences in data representing caused by variations in model architectures and training schemes. Moreover, adversarial perturbations tend to overfit target models during training, reducing their effectiveness across different models. To improve adversarial transferability, various strategies have been proposed. For example, Lin et al. [18, 34] propose data augmentation to increase data diversity. Xiong et al. [38, 4] propose ensemble-based attacks, which train perturbations against multiple surrogate models, based on the intuition that perturbations effective against multiple models are more transferable. However, this approach requires access to multiple models and incurs additional computational costs. Dong et al. [23, 13, 8] develop momentum strategies to stabilize update directions and prevent perturbations from getting stuck in poor local optima by accumulating historical gradients. A more comprehensive discussion of these methods can be seen in the survey [50].
II-B2 Unimodal and multi-modal attacks
Current methods primarily concentrate on unimodal cases, e.g., classification [31, 27, 13], where the objective is simply to disrupt the relationship between input data and its ground-truth label. In the real world, many scenarios involve multiple data modalities, e.g., image and text, leading to the development of multimodal models for tasks like image-text retrieval [51, 53], image captioning [6], and visual question answering [1]. To assess the robustness of multimodal models, multi-modal attack methods have been introduced, typically designed for specific tasks. For example, Zhang et al. [44, 52] learn adversarial perturabtions to attack image-text searching models by maximizing the neighbourhood relationships defined by a similarity matrix. Yin et al. [40] produce adversarial perturbations to mislead visual question answering models by preventing a substitute large language model from giving correct answers for each masked text. Since different tasks rely on distinct relationships between data modalities, attacks designed for one task often fail to transfer to others. This challenge has driven task-agnostic attack methods, including those for VLP models.
II-B3 Adversarial attack against VLP models
To address this limitation, Zhou et al. [55] explore the use of universal adversarial patches, though their approach is restricted to CLIP. Additionally, the adversarial patch is not strictly constrained in perturbation magnitude, making it more susceptible to detection. To further improve universality and transferability, Zhang et al. [47] propose learning universal adversarial perturbations by enhancing their local utility and incorporating ScMix data augmentation to increase diversity and reduce overfitting.
@@ II-B3 Adversarial attack against VLP models
Adversarial attacks against VLP models are still in their early stages, with only a few methods proposed. Compared to conventional multimodal attacks, attacking VLP models is more challenging due to the lack of access to downstream tasks and fine-tuning details. Existing methods primarily concentrate on the feature space and learn sample-wise perturbations by enlarging the feature distance between adversarial data and original data [42, 24, 43]. To further improve transferability, Lu et al. [24, 43] propose data augmentation to create more diverse image-text pairs, aiming to further exploit cross-modal interactions and mitigate overfitting. Despite the progress made, current approaches usually learn sample-wise perturbations, which are tailored to the specific characteristics and vulnerabilities of individual data samples. Consequently, when encountering new data, attackers have to learn tailored perturbations for each new data from scratch. This would incur huge computational costs and lead to less scalability for large-scale scenarios. To solve this, Zhou et al. [55] explore the use of universal adversarial perturbations, though their approach is restricted to CLIP. Zhang et al. [47] proposes to learn universal adversarial perturbations by enhancing their local utility and leveraging a ScMix data augmentation to promote data diversity to prevent overfitting. However, these methods focus solely on image-modality attacks and neglect the text modality, thereby limiting their effectiveness. Fang et al. [11] present an early attempt to generate both image and text attacks. Their method trains generators to produce adversarial examples by modeling data distributions, which risks overfitting to the source distribution. Moreover, it does not account for the characteristics of diverse downstream tasks and relies on an existing word library for text perturbation search, limiting its generality. In addition, the method generates adversarial text embeddings and then searches for semantically similar tokens in the corpus, creating a mismatch between embedding-level optimization and the final token-level perturbations.

### v2

@@ Abstract
To overcome this limitation, we propose Hierarchical Refinement Attack (HRA), a multimodal universal attack framework for VLP models. For the image modality, we refine the optimization path by leveraging a temporal hierarchy of historical and estimated future gradients to avoid local minima and stabilize universal perturbation learning. For the text modality, it hierarchically models textual importance by considering both intra- and inter-sentence contributions to identify globally influential words, which are then used as universal text perturbations. Extensive experiments across various downstream tasks, VLP models, and datasets, demonstrate the superior transferability of the proposed universal multimodal attacks.
@@
Vision-Language Pre-trained (VLP) models play a crucial role in bridging the gap between images and texts, enabling multimodal understanding, generation, and response [26, 13, 35]. Due to their promising performance and efficiency, VLP models have been widely applied to various vision-language tasks, ranging from multimodal retrieval [19] to image captioning [49]. As the deployment of VLP models grows, evaluating their robustness has become essential, particularly in high-stakes applications where reliability is critical. Uncovering model vulnerabilities would drive advancements in the development of more resilient and reliable neural networks. Adversarial attacks are a key strategy for probing model vulnerabilities and have been extensively studied [28, 45, 22, 2, 20, 43, 39]. These attacks typically involve crafting imperceptible perturbations that, when applied to input data, induce incorrect model predictions.
@@
Adversarial attacks are originally developed for unimodal settings, where only a single data modality is considered [28, 44, 22]. Recent work has explored attacks on specific multimodal tasks, such as image–text retrieval [40, 47, 48, 46] and visual question answering [36]. These approaches typically rely on task-specific losses or model outputs, limiting their applicability to task-agnostic VLP models. Current VLP attack methods focus on features that are independent of specific tasks and output, i.e., enlarging the feature distance between adversarial data and original data to break cross-modal alignments. However, current methods [38, 21, 39] generally customize adversarial perturbations for each sample based on individual characteristics. These sample-wise approaches have to learn adversarial perturbations for new data from scratch, resulting in significant computational overhead. This would significantly deter their applications from large-scale cases.
Zhou et al. [50, 42] make early attempts to learn universal adversarial attacks for the image modality. However, the method in [50] is restricted to specific models, and neither work exploits text-based attacks, which could further enhance adversarial effectiveness. Fang et al. [8] consider both image and text universal attacks, where the text-side attacks are implemented through word substitutions. These substitutions are obtained by first learning adversarial embeddings and then searching the corpus for words that are semantically aligned with these embeddings. However, this method requires access to a predefined word library, and the mismatch between the learned adversarial embeddings and their final token-level realizations can weaken the attack effectiveness. In addition, current methods are developped in the black-box settings, where the victim model is inaccessible and a surrogate model is leveraged as the target. However, different VLP models are pre-trained with diverse architectures, learning objectives, and datasets, and are further fine-tuned for specific downstream tasks [26, 13, 35]. These variations cause different VLP models to rely on distinct visual features or patterns. For instance, image–text retrieval primarily relies on global image–text similarity, while image captioning demands fine-grained visual recognition and semantic reasoning. Existing attacks are prone to overfitting to the surrogate model and failing to transfer to different downstream tasks, models, and data.
@@
Considering this, we propose Hierarchical Refinement Attack (HRA), which enhances the universal adversarial transferability to disrupt cross-modal alignments. Due to the fundamentally different natures of visual and textual data, HRA adopts modality-specific refinement strategies. Specifically, for images, which are continuous, we investigate the optimization trajectory, which is closely related to overfitting. Overfitting often occurs when the optimization process of gradient-based methods converges to local optima. To mitigate this problem, we propose future-aware momentum, which incorporates historical and predicted future gradients to regularize the current update and prevent the optimization from local optima. For text, which is discrete, we implement a simple yet effective adversarial text attack that leverages intra-sentence and inter-sentence importance measures to extract influential words as uniform adversarial perturbations. Extensive experiments demonstrate the effectiveness of the proposed method.
The main contributions of this work are:
A novel multi-modal attack method - HRA is proposed, which learns UAPs for both image and text modalities. The generated UAPs can be used for new data, tasks and models without retraining.
HRA hierarchically refines UAPs by leveraging past and future gradients to regularize the optimization trajectory, together with intra- and inter-sentence importance measures to enhance adversarial transferability.
@@
Thorough experiments are conducted on a wide range of target VLP models in multiple downstream tasks against various datasets. The results demonstrate that the proposed method effectively improves the transferability of universal multimodal attacks.
@@
The remainder of the paper is organized as follows: Section II provides a brief review of related work. Section III presents a detailed introduction to our methodology. In Section IV, we conduct extensive experiments, including attacks on various VLP models, datasets, and downstream tasks, as well as ablation studies and visualizations, to demonstrate the effectiveness of the proposed method. Finally, Section V concludes the paper and discusses current limitations and potential directions for future research.
@@ II-A Vision-Language Pre-training
Digital and realistic worlds encompass diverse data types, e.g., image, text, audio, video, and so on. To better understand the world, vision-language pre-training (VLP) has been developed to understand visual and language content, capture their relations and generate multimodal content. It harnesses vast amounts of unlabeled multi-modal data, using self-supervised learning techniques to gain representative feature learning ability [12, 26, 13, 35, 41]. VLP models exhibit outstanding generalizability and transferability, allowing adaptation to diverse datasets and downstream tasks through fine-tuning. Representative VLP methods include CLIP [26], BLIP [12], ALBEF [13], and TCL [35]. CLIP aligns visual-linguistic modalities by projecting images and texts into a unified feature space, using contrastive learning to maximize similarity for matched pairs and increase the distance for unmatched pairs. BLIP [12] enhances dataset quality by synthesizing captions for web images, employing three contrastive learning objectives to jointly pre-train the model, i.e., image-text contrastive learning, image-text matching, and image-conditioned language modelling. ALBEF [13] first aligns unimodal representations of pair images and texts, and then fuses them with cross-modal attention to obtain joint representations. TCL [35] utilizes contrastive learning to perform both inter- and intra-modal alignment while preserving mutual information between global and local representations in each modality.
@@ II-B Adversarial Attack
Real-world applications often involve uncertainty, randomness, and potential threats, making it crucial to ensure that developed models are both robust and secure. Adversarial attacks are one of the most effective methods for evaluating model robustness, which aims to induce incorrect predictions of target models by injecting imperceptible perturbations (i.e., adversarial perturbations) into input data [28, 40, 47, 45]. According to the scope, existing adversarial perturbations can be roughly categorized into sample-wise [22, 2], class-wise [1] and universal adversarial perturbations (UAPs) [23, 14, 17], which are designed for individual samples, specific classes, and all data, respectively. Generally, sample-wise and class-wise perturbations are more effective than universal perturbations as they can exploit vulnerabilities specific to each sample or class. However, they have to be generated from scratch for new samples or classes, limiting their scalability in large-scale applications. In contrast, universal perturbations are more efficient, as they can be reused across different data once learned. Nevertheless, they fail to leverage data-specific characteristics and are more prone to overfit to target models during training.
@@
Early adversarial attacks are developed in white-box settings, assuming full access to target models, tasks, and data [28, 40, 47]. In practice, such information is often unavailable, and generating perturbations for all possible targets is infeasible. This motivates black-box attacks, where target information is inaccessible. Existing approaches mainly fall into query-based attacks [24, 4, 51] and transfer-based attacks [18, 6, 32].
Query-based attacks infer adversarial perturbations by analyzing model feedback on perturbed inputs and typically estimate gradients via techniques like finite-difference approximations [4]. However, they still require access to the output of target models, limiting practicality. Transfer-based attacks alleviate this constraint by leveraging substitute models, but their effectiveness and transferability is often compromised due to overfitting to substitute models. To handle this, several strategies have been proposed. Data augmentation is employed to increase input diversity [15, 31], while ensemble-based attacks optimize perturbations against multiple surrogate models to improve generalization [34, 3]. However, ensemble methods require access to multiple models and incur additional computational overhead. Momentum-based attacks stabilize update directions by accumulating historical gradients, thereby reducing convergence to poor local optima and improving transferability [20, 10, 5].
II-B2 Adversarial attack against VLP models
@@
Adversarial attacks against VLP models are still in their early stages, with only a few methods proposed. Compared to conventional multimodal attacks, attacking VLP models is more challenging due to the lack of access to downstream tasks, models and fine-tuning details. Existing methods primarily concentrate on the feature space and learn sample-wise perturbations by enlarging the feature distance between adversarial data and original data [38, 21, 39]. To further improve transferability, Lu et al. [21, 39] propose data augmentation to create more diverse image-text pairs, aiming to further exploit cross-modal interactions and mitigate overfitting. Despite the progress made, current approaches usually learn sample-wise perturbations, which are tailored to the specific characteristics and vulnerabilities of individual data samples. Consequently, attackers have to learn tailored perturbations for each new data from scratch. This would incur huge computational costs and lead to less scalability for large-scale scenarios. To solve this, Zhou et al. [50] explore the use of universal adversarial perturbations, though their approach is restricted to CLIP. Zhang et al. [42] proposes to learn universal adversarial perturbations by enhancing their local utility and leveraging a ScMix data augmentation to promote data diversity to prevent overfitting. However, these methods focus solely on image-modality attacks and neglect the text modality, thereby limiting their effectiveness. Fang et al. [8] present an early attempt to generate both image and text attacks. Their method trains generators to produce adversarial examples by modeling data distributions, which risks overfitting to the source distribution. In addition, the method generates adversarial text embeddings and then searches for semantically similar tokens in the corpus, creating a mismatch between embedding-level optimization and the final token-level perturbations.